Solve Linear systems by multiplying first.solve the linear system using elimination.
7x+3y=-12
2x+5y=38

Answer by mananth(16946)

7 x + 3 y = -12 .............1
2 x + 5 y = 38 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 3
-35 x -15 y = 60
6 x 15 y = 114
Add the two equations
-29 x = 174
/ -29
x = -6
plug value of x in (1)
7 x + 3 y = -12
-42 + 3 y = -12
3 y = -12 + 42
3 y = 30
y = 10